The following equation is 2H2 + O2---2H2O



If 5.0 grams of H2 reacts completely with O2, how many grams of H2O will be produced?

1
Expert's answer
2022-04-16T04:09:26-0400

2H2 + O2 = 2H2O

According to the equation, n (H2O) = n (H2).

n = m / M

M (H2O) = 18 g/mol

M (H2) = 2 g/mol

n (H2) = 5.0 / 2 = 2.5 mol

n (H2O) = n (H2) = 2.5 mol

m (H2O) = n x M = 2.5 x 18 = 45 g